# FlorDB: Multiversion Hindsight Logging for Continuous Training

### 06 Feb 2024 v2

@@ Abstract.
Production Machine Learning involves continuous training: hosting multiple versions of models over time, often with many model versions running at once. When model performance does not meet expectations, Machine Learning Engineers (MLEs) debug issues by exploring and analyzing numerous prior versions of code and training data to identify root causes and mitigate problems. Traditional debugging and logging tools often fall short in managing this experimental, multi-version context. FlorDB introduces Multiversion Hindsight Logging, which allows engineers to use the most recent version’s logging statements to query past versions, even when older versions logged different data. Log statement propagation enables consistent injection of logging statements into past code versions, regardless of changes to the codebase. Once log statements are propagated across code versions, the remaining challenge in Multiversion Hindsight Logging is to efficiently replay the new log statements based on checkpoints from previous runs. Finally, a coherent user experience is required to help MLEs debug across all versions of code and data. To this end, FlorDB presents a unified relational model for efficient handling of historical queries, offering a comprehensive view of the log history to simplify the exploration of past code iterations. We present a performance evaluation on diverse benchmarks confirming its scalability and the ability to deliver real-time query responses, leveraging query-based filtering and checkpoint-based parallelism for efficient replay.
@@ 1. Introduction
Model development for machine learning (ML) differs from traditional software engineering, focusing on experimental iteration rather than conventional debugging. The primary goal is to boost the predictive performance of models, leading developers to adopt empirical methods involving thorough logging and continuous updates to code and data (Xin et al., 2021). A recent study highlighted the importance of high-speed experimentation in this field: one participant stated, “the most important thing to do is achieve scary high experimentation velocity” (Shankar et al., 2024). This approach, however, generates a significant data management challenge, as engineers must handle numerous iterations of code, datasets, and logs.
For this aggressively agile approach to be successful, machine learning engineers (MLEs) must be able to retroactively examine past results to guide ongoing experiments (Garcia et al., 2018). One method to enable model developers to “query the past” is by generating and maintaining comprehensive experiment logs, but this is burdensome and unrealistic. Ensuring comprehensive logs requires (a) foresight to add logging statements to the code for all relevant metrics, without knowing which ones are relevant in advance, and (b) data management discipline to manage the corpus of log files (and requisite metadata) that ensues.
@@ 1.1. Multiversion Hindsight Logging
Hindsight logging (Garcia et al., 2020) is a lightweight record-replay technique that allows MLEs to add logging statements to a long-running Python program after it executes, and then incrementally replay parts of the program very quickly to generate the log outputs that would have emerged had the statements been in the code originally. This technique encourages MLEs to optimistically and flexibly operate with a minimal initial logging scheme, typically restricted to loss and accuracy metrics during the training phase. Additional logging is added retroactively, when evidence of issues arises (e.g., in deployment) and further key information is required for debugging.
@@ 1.2. Contributions
FlorDB, a Multiversion Hindsight Logging system, provides a SQL-like query interface for ease of use, treating every version run as a set of rows, and each logging statement as a “virtual column”. FlorDB includes automatic version control through Git and is designed to interoperate with multiple ML experiment management tools. In addressing the inherent challenges of implementing Multiversion Hindsight Logging, FlorDB makes the following key contributions:
@@
    Multiversion Hindsight Logging as Acquisitional Query Processing: FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005) by seamlessly integrating data acquisition into query execution through experiment replay. This framework differs from traditional systems that query static datasets. Selected experimental versions undergo sequential processing: logging statements are propagated to generate required data, followed by the partial or parallel execution of a modified training script.
@@
    Highly Accurate Time Estimation for Replay Query Refinement: FlorDB incorporates highly accurate cost prediction for estimating the runtime of on-demand replay queries. By leveraging runtime profiling statistics collected during the initial record phase, FlorDB provides highly accurate (error $\leq 5\%$) replay cost estimates. This precision enables developers to refine their queries before being bogged down by unnecessarily lengthy execution times.
@@ 2. Scenario: Catastrophic Forgetting
Figure 1. Average training losses and F1-scores for Alice’s object detection model over the last 6 months. The model undergoes continuous training, with batches of labeled data added approximately twice a month. These batch dumps result in temporary fluctuations in the loss. The overall trend is stable.
@@
Catastrophic forgetting is a pressing issue in machine learning and is especially problematic in production settings (Shankar et al., 2024). This phenomenon occurs when a model, upon being fine-tuned for a new task, loses its ability to perform well on a previously learned task. This creates a balancing act between updating the model for new data or tasks and maintaining its performance on older tasks.
@@ 2.1. Object Detection in Autonomous Vehicles
However, Alice encounters an unexpected problem. A colleague, Bob, informs her of persistent failures in pedestrian detection. Upon examining a representative set of images (refer to the bottom row of Figure 2), she verifies the issue.
@@ 3. User Experience & API
    Reads a value from the command line if provided; otherwise, uses the default value. During replay, retrieves historical values. Good for setting configuration parameters and allowing for quick adjustments via the command line.
@@
    Acts as a Python context manager that defines the set of objects, such as model or optimizer, to be checkpointed adaptively — based on the size of the checkpoint and the loop iteration time (Garcia et al., 2020) — at flor.loop iteration boundaries.
@@ 3.1. Running Experiments
Alice executes her model training via the following command that starts her train.py script with specific hyper-parameters:
@@ 3.1.1. Args & Logs
In the context of running experiments, Alice utilizes the Flor API to enhance configurability and traceability. The flor.arg function is instrumental in parameterizing her experiments, allowing her to define hyper-parameters such as hidden_size, learning_rate, batch_size, and seed either via the command-line or by utilizing default values specified within the script (lines 5-8 in 3). This ensures adaptability and reproducibility of her experiment’s configuration. In addition, flor.log is employed to record the experiment’s loss metrics and associated metadata at each optimizer step (line 20 in 3). The logging of the loss captures the dynamics of the model’s performance, enabling Alice to monitor the training process.
@@ 3.1.2. Checkpointing on Loop Boundaries
Alice’s code defines the set of objects to be checkpointed using flor.checkpointing (line 10); within that context, Alice calls flor.loop (lines 12-13) to use the checkpoints and control record-replay. After each iteration of the outermost loop, which here corresponds to the completion of an “epoch,” FlorDB assesses whether to checkpoint. This decision is based on a balance between the checkpoint’s size and the iteration’s duration, adhering to a strategy from prior work (Garcia et al., 2020). Checkpoints are created adaptively, approximately once per epoch.
@@ 3.2. Experiment Analysis and Hindsight Logging
Upon initiating the replay command, Alice is shown a schedule of experiments slated for replay and must confirm to proceed. After receiving confirmation, Flor restores historical configurations using flor.arg. Notably, during this replay, Flor bypasses the main loop (lines 12-22 in 3) to directly load the model’s final state, as F1-scores are computed post-training (line 25 in 3).
Alice views the updated results with the following command:
@@
This command now generates a table including the f1_ped metric (green text in 4) for entries dated June 23, 2023, or later. This new data reveals a previously unnoticed performance regression in pedestrian classifications, as Bob had indicated (tstamp $\geq$ 08-27).
To further refine the analysis and calibrate alert thresholds, Alice updates the validation function to add a new flor.log statement as part of an assertion that pedestrian F1-scores never fall below some threshold (line 21 in 3). For retroactive application, she again employs FlorDB’s replay feature:
@@ 4. System Architecture
FlorDB maintains a cross-referenced relational database that interconnects code versions in git, large files in your preferred storage solution (e.g. an S3 bucket), and metadata stored in the database. The design of FlorDB aims to provide a unified platform for the management of code, checkpoints, logs, and their evolution over time (5). This section outlines the mechanisms FlorDB uses for executing experiments and replaying them for hindsight logging, highlighting how it stores, manages and tracks the data and metadata of model training.
@@ 4.1. Storage and Data Layout
  • Git Repository: Central to version control, it captures the state of the working directory and JSON log files post-execution. This facilitates a comprehensive tracking of code versions and more transparent sharing of execution metadata, enhancing experiment reproducibility and analysis.
  • Relational Database: Following execution and logging to JSON, FlorDB unpacks and normalizes data from the JSON logs to populate the relational database. This process transforms the semi-structured JSON data into a structured format, suitable for efficient querying and analysis.
@@
This structured approach to storage ensures that FlorDB is well-equipped to manage the complexities of code versioning, data checkpointing, and metadata storage.
@@ 4.2. Relational Schema
  • projid: Basename of the cwd or Git repo.
@@
The “unpack” process involves converting logging objects from JSON format into the relational structure of the database. Each JSON log record is parsed. The data is then mapped to the relevant columns in both the logs and loops tables: A null loop indicating a top-level log entry, the log record is mapped solely within the logs table; nested loop objects encapsulate the hierarchical structure of an experiment’s iterations, and the corresponding log record is reflected in both logs and loops tables.
@@
4.3. Flor DataFrame
The flor.dataframe presents the content of the database in a single-table view (7). The operations involved join and transform the relational schema (6) into the final view. Initially, the code joins the logs table with additional dimensions of data, such as the loops table, which are essential for capturing the full context of each experimental run. This is followed by a series of transformations that pivot the combined data, turning what were once values into column headers, to create a wide-format table where each unique loop and log entry becomes a distinct column in the DataFrame (e.g. 4).
@@ 4.3.1. Domain Mapping
We use the flor.dataframe as the default Flor view because of its easy to understand application-level semantics. Specifically:
@@
The ML developer can envision the flor.dataframe as a universal relational view over the columns they have selected, where the rows represent individual experiment runs or, in the presence of loops, iterations within those runs. The columns in the flor.dataframe view are potentially infinite; they can be defined post-hoc and subsequently populated using replay from a checkpoint. This fluidity in defining and back-filling data exemplifies the relational model’s capacity to adapt to an experimenter’s evolving analytical needs.
@@ 5.2. Empirical Cost Prediction & Plan Execution
Machine Learning (ML) training scripts can be long-running, and even selective replay can be time-consuming. Therefore, ML engineers need to know how long a replay will take, to decide whether further query refinement is needed. Although query cost estimation is a notoriously hard problem (Selinger et al., 1979; Yang et al., [n.d.]), we can accurately estimate costs for replay executions. This is because FlorDB profiles runtime information during the record phase. When initiating a Flor replay, these runtime statistics are queried using the flor.dataframe API call. This allows us to accurately estimate the time required for replay, enabling users to tailor their queries for efficiency. We categorize the cost estimation into four levels: s
@@
    • Cost Estimate: sum(prefix)
    • Subprocess: python train.py --replay_flor 0=-1
@@
    • Cost Estimate: sum(prefix + ckpt_load + suffix)
    • Subprocess: python train.py --replay_flor 0=0
@@
    • Estimate: sum(prefix + N*(ckpt_load + val_margin))
      where N is the number of epochs.
    • Subprocess: python train.py --replay_flor 0=1 1=0
@@
    Range Scan:Runs the training loop in-depth (lines 12-22 in 3) over a selected range to capture nested logging statements — used for the most granular analyses (e.g. of gradients). This level can be used individually or as part of a full replay in parallel, where each GPU is assigned a range of work.
    • Estimate: sum(prefix + ckpt_load + (k/N)*outer)
      where outer is the profiled time to run all N epochs, and k is the length of the selected range.
    • Subprocess: Let SEG be a list of int values denoting a contiguous range of epochs $i..j$ s.t. $0<i<j\leq N$; then,
      python train.py --replay_flor 0=\$SEG 1=1
@@ 5.2.1. Plan Execution
Once the replay’s depth level has been determined by a static analysis of the training script, and the end-user has accepted the replay plan, a replay subprocess is invoked for each version in the plan, with --replay_flor arguments formatted as strings and flags guiding Flor’s operations. As shown in the list above, Flor replay arguments can be used to break out of the execution in a prefix scan upon encountering the first flor.loop (--replay_flor 0=-1 where 00 is the index of the first flor.loop); or set to skip the first flor.loop loading its end checkpoint as part of a suffix scan (--replay_flor 0=0); or set to step into the the first flor.loop (--replay_flor 0=1), and so on. As shown above, validation scans and range scans control access to the outermost loop (index 00 or "epoch") and the nested loop (index $1$ or "step"), a key difference being that in validation scan the nested loop is skipped (1=0) but not in the range scan (1=1).
@@ 5.3. Logging Statement Propagation
After selecting the versions and log variables for replay, but before the execution of individual Flor replay subprocesses, FlorDB adapts the code from the Git repository to incorporate new log statements. This involves propagating log statements from a newer version $Y$ to an older version $X$ ($X<Y$). The core challenge here is one of code-block alignment: identifying the most appropriate line in version $X$ to isert a logging statement that was originally in line $L$ of version $Y$. Code alignment is a fundamental task in software engineering, essential for understanding and managing changes in source files (Horwitz, 1990; Ladd, 1994). It allows developers to track modifications and maintain consistency across different versions of a file. In this context, GumTree emerges as a state-of-the-art technique for code block alignment.
@@
Code alignment, crucial for tracking and managing changes, uses GumTree (Falleri et al., 2014) for aligning code blocks, aiding in the accurate insertion of log statements across versions. The flor.loop() function establishes stable anchor points, with main and nested training loops serving as consistent markers across versions. This consistency, enhanced by GumTree’s alignment capabilities, ensures accurate log statement propagation. While GumTree is capable of aligning code blocks without Flor API anchors, the presence of these static loop identifiers all but ensures its efficiency and accuracy: in all instances of logging statement propagation evaluated in 6, this combined approach was successful.
To assess the effectiveness of these techniques, we conducted an evaluation focusing on accuracy and comparing GumTree with the Myers algorith (Myers, 1986; Falleri et al., 2014). The evaluation included synthetic benchmarks, randomly mutated PyTorch programs, and ecological benchmarks from GitHub. Our findings, detailed in a technical report (Dandamudi, 2021), show GumTree’s superior performance, especially in handling code refactoring and variable renaming.

### 02 Mar 2024 v3

@@ Abstract.
Production Machine Learning is a data-intensive process, involving continuous retraining of multiple versions of models, many of which may be running in production at once. When model performance does not meet expectations, Machine Learning Engineers must explore and analyze numerous versions of code, logs and training data to identify root causes and mitigate problems. Traditional software engineering tools fall short in this data-rich context. FlorDB introduces multiversion hindsight logging, a form of acquisitional query processing that allows engineers to use the most recent version’s logging statements to query past versions’ logs, even when older versions logged different data. FlorDB provides a replay query interface with accurate cost estimates to help the end-user refine their queries. Once a replay query plan is confirmed, logging statements are propagated across code versions, and the modified training scripts are replayed based on checkpoints from previous runs. Finally, FlorDB presents a unified relational view of log history across versions, making it easy to explore behavior across past code iterations. We present a performance evaluation on diverse benchmarks confirming scalability and the ability to deliver real-time query responses, leveraging query-based filtering and checkpoint-based parallelism for efficient replay.
@@ 1. Introduction
Model development for machine learning (ML) is an iterative, experimental and data-intensive process that differs from traditional software engineering. The primary goal in model development is to boost predictive performance, leading developers to adopt empirical methods involving thorough logging and continuous updates to code and training data (Xin et al., 2021). A recent study highlighted the importance of high-speed experimentation in this field: as one participant stated, “the most important thing to do is achieve scary high experimentation velocity” (Shankar et al., 2024). This approach, however, generates a significant data management challenge for engineers, who must handle numerous iterations of code, datasets, and logs.
For this aggressively agile approach to be successful, machine learning engineers (MLEs) must be able to retroactively examine rich and diverse data from past runs to guide ongoing experiments (Garcia et al., 2018). One method to enable model developers to “query the past” is by generating and maintaining comprehensive experiment logs, but this is burdensome and unrealistic. Ensuring comprehensive logs requires (a) foresight to add logging statements to the code for all relevant metrics, without knowing which ones are relevant in advance, and (b) data management discipline for the many log files, code versions, training data and requisite metadata that ensues.
@@ 1.1. Multiversion Hindsight Logging
Hindsight logging (Garcia et al., 2020) is a lightweight record-replay technique that allows MLEs to add logging statements to a long-running Python program after it executes, and then incrementally replay parts of the program very quickly to generate the log outputs that would have emerged had the statements been in the code originally. This technique encourages MLEs to operate optimistically and flexibly with a minimal initial logging scheme, typically restricted to loss and accuracy metrics during the training phase. Additional logging is added retroactively, when evidence of issues arises (e.g., in deployment) and further key information is required for debugging.
@@ 1.2. Contributions
FlorDB, a Multiversion Hindsight Logging system, provides a tabular query interface, treating every run of a version as a set of rows, and each logging statement as a “virtual column”. FlorDB includes automatic version control through Git and is designed to interoperate with multiple ML experiment management tools. In addressing the inherent challenges of implementing Multiversion Hindsight Logging, FlorDB makes the following key contributions:
@@
    Multiversion Hindsight Logging as Acquisitional Query Processing: FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005) by seamlessly integrating data acquisition into query execution through experiment replay. This framework differs from traditional systems that query static datasets. Selected experimental versions undergo sequential processing: logging statements are propagated to generate required data, followed by the partial or parallel replay of modified training scripts.
@@
    Accurate Time Estimation for Query Refinement: FlorDB incorporates highly accurate cost prediction for estimating the runtime of on-demand replay queries. By leveraging runtime profiling statistics collected during the initial record phase, FlorDB provides highly accurate (error $\leq 5\%$) replay cost estimates. This precision enables developers to refine their queries before being bogged down by unnecessarily lengthy execution times.
@@ 2. Scenario: Catastrophic Forgetting
Figure 1. Average training losses and F1-scores for Alice’s object detection model over the last 6 months. The model undergoes continuous training, with batches of labeled data added approximately twice a month. These batch dumps result in temporary fluctuations in the loss. F1-round is the F1-score for roundabouts; F1-score is the global F1-score.
@@
For readers who are not MLEs, we motivate FlorDB with an example of catastrophic forgetting, a pressing ML issue that is especially problematic in production settings (Shankar et al., 2024). This phenomenon occurs when a model, upon being fine-tuned for a new task, loses its ability to perform well on a previously learned task. This creates a balancing act between updating the model (for new data or tasks) and maintaining its performance (for older tasks).
@@ 2.1. Object Detection in Autonomous Vehicles
However, Alice encounters an unexpected problem. A colleague, Bob, informs her of persistent failures in pedestrian detection that have arisen in her new model versions. Upon examining a representative set of images (refer to the bottom row of Figure 2), she verifies the issue.
@@ 3. User Experience & API
    Reads a value from the command line if provided; else uses the default value. During replay, retrieves logged values from history. Good for setting configuration parameters and allowing for quick adjustments via the command line.
@@
    Scopes the set of objects to be checkpointed, such as a model or optimizer, via a Python context manager. Checkpointing is done adaptively at flor.loop iteration boundaries, based on the checkpoint size and the loop iteration time (Garcia et al., 2020).
@@ 3.1. Running Experiments
Returning to our scenario, Alice executes her model training via the following command that starts her train.py script with specific hyper-parameters:
@@ 3.1.1. Args & Logs
In the context of running experiments, Alice utilizes the Flor API to enhance configurability and traceability. The flor.arg function (lines 5-8 in 3) is instrumental in parameterizing her experiments, allowing her to define hyper-parameters such as hidden_size, learning_rate, batch_size, and seed either via the command-line or by utilizing default values specified within the script. This ensures adaptability and reproducibility of her experiment’s configuration. In addition, flor.log is employed (line 20 in 3) to record the experiment’s loss metrics and associated metadata at each optimizer step. The logging of the loss captures the dynamics of the model’s performance, enabling Alice to monitor the training process.
@@ 3.1.2. Checkpointing on Loop Boundaries
Alice’s code defines the set of objects to be checkpointed using flor.checkpointing (line 11); within that context, Alice calls flor.loop (lines 12-13) to use the checkpoints and control record-replay. After each iteration of the outermost loop, which here corresponds to the completion of an “epoch,” FlorDB assesses whether to checkpoint. This decision is based on a balance between the checkpoint’s size and the iteration’s duration, adhering to a strategy from prior work (Garcia et al., 2020). Checkpoints are created adaptively, approximately once per epoch.
@@ 3.2. Experiment Analysis and Hindsight Logging
Upon initiating the replay command, Alice is shown a schedule of experiments slated for replay and must confirm to proceed. After receiving confirmation, Flor prepares to replay the selected prior runs: it restores their versions from github, adds the new f1_ped logging statement to each, and executes all versions using the configuration variable values originally logged via flor.arg. Notably, during this replay, Flor’s hindsight logging functionality can bypass the main loop (lines 12-22 in 3) to directly load the model’s final state, since the f1_ped statement is computed post-training (line 25 in 3).
Once replay is finished, Alice views the updated results with the following command:
@@
This command now generates a table including the f1_ped metric (green text in 4) for entries dated June 23, 2023, or later. This new data reveals a previously unnoticed performance regression in pedestrian classifications, as Bob had indicated (f1_ped values in the range tstamp $\geq$ 08-27).
To further refine the analysis and calibrate alert thresholds, Alice updates the validation function to add a new flor.log statement as part of an assertion that pedestrian F1-scores never fall below some threshold. For retroactive application, she again employs FlorDB’s replay feature:
@@ 4. System Architecture
The design of FlorDB aims to provide a unified platform for the management of code, checkpoints, logs, and their evolution over time (5). FlorDB maintains a cross-referenced relational database that interconnects code versions in git, large files in your preferred storage solution (e.g. an S3 bucket), and metadata stored in the database. This section outlines the mechanisms FlorDB uses for executing experiments and replaying them for hindsight logging, highlighting how it stores, manages and tracks the data and metadata of model training.
@@ 4.1. Storage and Data Layout
  • Git Repository: For code version control, FlorDB uses git to capture the state of the working directory and JSON log files post-execution. FlorDB and the user commit to the same Git repository, but FlorDB auto-commits to a dedicated branch for added safety — the user may treat that branch as any other branch. This facilitates a comprehensive tracking of code versions and more transparent sharing of execution metadata, enhancing experiment reproducibility and analysis.
  • Relational Database: Following execution and logging to JSON, FlorDB unpacks and normalizes data from the JSON logs to populate a relational database (which is pluggable; FlorDB uses SQLite by default). This process transforms the semi-structured JSON data into a structured format, suitable for efficient querying and analysis.
@@
This heterogeneous approach to storage ensures that FlorDB is well-equipped to manage the complexities of code versioning, data checkpointing, and metadata storage.
4.1.1. Serialization Layer
The API utilizes a serialization layer to prepare objects for storage. This layer differentiates between object types — PyTorch objects are serialized using native PyTorch functionality, while cloudpickle provides a general-purpose serialization solution. For in-depth information on how fork and copy-on-write are used for background serialization, the reader is referred to a technical report (Liu, 2020).
@@ 4.2. Relational Schema
  • projid: Basename of the root-level working directory.
@@
The “unpack” process involves converting logging objects from JSON format into the relational structure of the database. Each JSON log record is parsed. The data is then mapped to the relevant columns in both the logs and loops tables. In the case of a null loop indicating a top-level log entry, the log record is mapped solely within the logs table; in cases of nested loops, the JSON object encapsulates the hierarchical structure of an experiment’s iterations, and the corresponding log record is reflected in both logs and loops tables.
@@
4.3. Pivoted Views
The flor.dataframe presents the content of the database in a single-table view (7). This view is generated from the relational schema (6) by joining the logs table with additional dimensions of data, such as the loops table, which are essential for capturing the full context of each experimental run. This is followed by a series of transformations that pivot the combined data, turning relational values (e.g., logs.value_name, loops.loop_name) into column headers, to create a wide-format table where each unique loop and log entry name becomes a distinct column in the DataFrame (e.g. 4).
@@ 4.3.1. Domain Mapping
We use the flor.dataframe as the default Flor view because of its easy-to-understand application-level semantics. Specifically:
@@
The ML developer can envision the flor.dataframe as a universal relational view over the columns they have selected, where the rows represent individual experiment runs or, in the presence of loops, iterations within those runs. The columns in the flor.dataframe view are potentially infinite; they can be defined post-hoc and subsequently populated using replay from a checkpoint. This fluidity in defining and back-filling columns and data allows a relational model to adapt to an experimenter’s evolving analytical needs.
@@ 5.2. Empirical Cost Prediction & Plan Execution
Figure 8. Time estimation error per replay operator; micro-benchmarks correspond to a 2-layer neural network. [^3] See train.py in https://github.com/ucbepic/ml_tutorial
Machine Learning (ML) training scripts can be long-running, and even selective replay can be time-consuming. Therefore, ML engineers need to know how long a replay will take, to decide whether further query refinement is needed. Traditional query cost estimation is a notoriously hard problem, often exhibiting orders of magnitude errors in estimation (Leis et al., 2015). By contrast, we can estimate query runtime for replay executions extremely accurately (3). This is because FlorDB profiles runtime information during the record phase. When initiating a Flor replay, these runtime statistics are queried using the flor.dataframe API call. This allows us to reliably predict the time required for replay, enabling users to tailor their queries for efficiency. We categorize the cost estimation into four levels:
@@
    Range Scan: Runs the training loop in depth (lines 12-22 in 3) over a selected range to capture nested logging statements — used for the most granular analyses (e.g. of gradients). A range scan running from epoch 0 to $N$ is considered a full scan. To achieve replay parallelism, FlorDB executes range scans over non-overlapping intervals.
@@ 5.2.1. Plan Execution
Once the replay’s depth level has been determined by a static analysis of the training script, and the end-user has accepted the replay plan, a replay subprocess is invoked for each version in the plan, with CLI flags and arguments formatted as strings parameterizing the replay’s depth level. Flor replay arguments can be used to break out of the execution in a prefix scan upon encountering the first flor.loop or set to skip the first flor.loop loading its end checkpoint as part of a suffix scan; or set to step into the the first flor.loop as part of a validation scan, and so on. As shown above, validation scans and range scans control access to the outermost loop, and the nested loop, a key difference being that in validation scan the nested loop is skipped but not in the range scan.
@@ 5.3. Logging Statement Propagation
After selecting the versions and log variables for replay, but before the execution of individual Flor replay subprocesses, FlorDB adapts the code from the git repository to incorporate new log statements. This involves propagating log statements from a newer version $Y$ to an older version $X$ ($X<Y$). The core challenge here is one of code-block alignment: we must identify the most appropriate line in version $X$ to insert a logging statement that was originally in line $L$ of version $Y$. Code alignment is a fundamental task in software engineering, essential for understanding and managing changes in source files (Horwitz, 1990; Ladd, 1994). It allows developers to track modifications and maintain consistency across different versions of a file. In this context, GumTree emerges as a state-of-the-art technique for code block alignment.
@@
FlorDB’s code alignment uses GumTree (Falleri et al., 2014) for aligning code blocks, aiding in the accurate insertion of log statements across versions. The flor.loop() function establishes stable anchor points, with main and nested training loops typically serving as consistent markers across versions. This consistency, enhanced by GumTree’s alignment capabilities, ensures accurate log statement propagation. While GumTree is capable of aligning code blocks without Flor API anchors, the presence of these static loop identifiers all but ensures its efficiency and accuracy: in all instances of logging statement propagation evaluated in 6, this combined approach was successful.
To assess the effectiveness of these techniques, we conducted an evaluation focusing on accuracy and comparing GumTree with the Myers algorithm (Myers, 1986; Falleri et al., 2014). The evaluation included synthetic benchmarks, randomly mutated PyTorch programs, and ecological benchmarks from GitHub. Our findings, detailed in a technical report (Dandamudi, 2021), show GumTree’s superior performance, especially in handling code refactoring and variable renaming.